OpenStudy (anonymous):

whats the name of this equation? x equals negative b plus or minus the square root of b squared minus 4ac all over two a.

OpenStudy (anonymous):

the general quadratic equation

OpenStudy (anonymous):

thanks i couldnt remember

OpenStudy (kinggeorge):

I would call it the quadratic formula. The general quadratic equation is generally written \(ax^2+bx+c\)

OpenStudy (anonymous):

call it what you want, they're the same equation :P \[ax^2 + bx + c = 0\]\[a(x+ \frac{b}{2a})^2 +c -\frac{b^2}{4a} = 0\] \[a(x+ \frac{b}{2a})^2 = \frac{b^2}{4a} -c\] \[(x+ \frac{b}{2a})^2 = \frac{b^2 -4ac}{4a^2}\] \[x+ \frac{b}{2a} = \pm \sqrt{\frac{b^2 -4ac}{4a^2}}\] \[x+ \frac{b}{2a} = \frac{ \pm \sqrt{b^2 -4ac}}{2a}\] \[x = \frac{-b \pm \sqrt{b^2 -4ac}}{2a}\]

OpenStudy (kinggeorge):

As long as the quadratic equation \(ax^2+bx+c\) equals 0. In general, that's not true. However, any quadratic equation can be made to equal 0 with ease.

OpenStudy (anonymous):

i don't really see the distinction.. but it doesnt really matter :D names are just names after all